OpenStudy (anonymous):

determine absolute or conditional convergence: sum(k=1 -> infinity) [(-1)^(k+1)(10^k)]/(k!)

OpenStudy (anonymous):

have u tried ratio test?

OpenStudy (anonymous):

well it's an alternating series. i think i'm supposed to use the alternating series test

OpenStudy (anonymous):

of course alternating series test will work :)

OpenStudy (anonymous):

i'm not sure how i'm supposed to do it because of the k+1 on the -1

OpenStudy (anonymous):

u just need to drop \[(-1)^{k+1}\]and then work on\[\frac{10^k}{k!}\]

OpenStudy (anonymous):

this is a good sourse http://tutorial.math.lamar.edu/Classes/CalcII/AlternatingSeries.aspx

OpenStudy (anonymous):

what is ur difficulty with that??

OpenStudy (anonymous):

ok.. now i get that.. so how should i take the limit of 10^k/k! ?

OpenStudy (anonymous):

oh sorry i was out ;) well k! is very greater than 10^k when k increses

OpenStudy (anonymous):

*increases

OpenStudy (anonymous):

so would that mean that the limit would go to zero?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how to tell weather it converges absolutely or conditionally?

OpenStudy (anonymous):

emm...u should apply Absolute Convergence test

OpenStudy (anonymous):

sorry.. i kinda don't understand how to do that here... could you help me out??

OpenStudy (anonymous):

ok u got the first part right? alternating series test?

OpenStudy (anonymous):

yes, lim(k>infinity)[10^k/k!] = 0

OpenStudy (anonymous):

therefore it converges

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

\[\lim_{k \rightarrow \infty}a_k=0\] Is a REQUIREMENT for a series to converge but it doesn't not IMPLY that a series converges.

OpenStudy (anonymous):

it does not*

OpenStudy (anonymous):

Even if it is an alternating series do the ratio test, the ratio test works well for factorials. Also, if that converges then you know it absolutely converges (because you take the absolute value so the (-1)^k+1 goes away) and if it absolutely converges then you know it conditionally converges.